OpenStudy (anonymous):

Two boats depart from a port located at (–10, 0) in a coordinate system measured in kilometers, and they travel in a positive x-direction. The first boat follows a path that can be modeled by a quadratic function with a vertex at (0, 5), and the second boat follows a path that can be modeled by a linear function and passes through the point (10, 4). At what point, besides the common starting location of the port, do the paths of the two boats cross?

OpenStudy (anonymous):

Possible Answers... (–6, 0.8) (–6, 3.2) (6, 3.2) (6, 0.8)

OpenStudy (anonymous):

:( sorry, those diamond question marks are both negative signs.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

I can see how to get the equation of boat 2 - the linear path. Maybe we need a graph.

Directrix (directrix):

Is this the same question: Two boats depart from a port located at (–10, 0) in a coordinate system measured in kilometers, and they travel in a positive x-direction. The first boat follows a path that can be modeled by a quadratic function with a vertex at (0, 5), and the second boat follows a path that can be modeled by a linear function and passes through the point (10, 4). At what point, besides the common starting location of the port, do the paths of the two boats cross? http://www.softmath.com/algebra-word-problems/show.php?id=20795

Directrix (directrix):

Do we agree with these imported statements regarding Boat 1? First boat the equation of a quadratic with vertex in (h, k) is given by y = a(x – h)2 + k in our case the vertex is (0, 5) that is h = 0 and k = 5 y = a(x - 0)^2 + 5 y = ax^2 + 5 also the line goes trough the point (-10, 0) 0 = a(-10)^2 + 5 a = -5/100 a = - 1/20 the path of the first boat is described by the equation y = (-1/20)x^2 + 5

Directrix (directrix):

Second boat the equation of the linear function through the points (-10, 0) and (10, 4) has a slope equal to: m = (4 - 0)/(10 - (-10)) = 4/20 = 1/5 y - y1 = m(x - x1) y - 0 = (1/5)(x - (-10) y = (1/5)x + 10/5 y = (1/5)x + 2 the path of the second boat is described by the equation y = (1/5)x + 2

Directrix (directrix):

At what point, besides the common starting location of the port, do the paths of the two boats cross? we fill find this by solving the system y = (-1/20)x^2 + 5 y = (1/5)x + 2 that is by solving the equation (-1/20)x^2 + 5 = (1/5)x + 2 we find the 2 roots x1 = 6 y1 = (1/5)6 + 2 = 6/5 +2 = 3.2 (6, 3.2) x2 = -10 y2 =(1/5)(-10) + 2 = 0 (-10, 0)

Directrix (directrix):

All of the work posted above is that of @Algebrator at http://www.softmath.com/algebra-word-problems/show.php?id=20795

OpenStudy (mathmath333):

https://www.desmos.com/calculator/w45arwv7eq

OpenStudy (anonymous):

Ok, so (6, 3.2) is answer, because, when graphed, the equations intersect at that point(and another, but that one wouldn't make sense).

OpenStudy (mathmath333):

they meet two times 1.) when they start at the staring point 2.) when they intersect

OpenStudy (anonymous):

They made it clear they didn't want #1. That is why I said it didn't make sense.

Directrix (directrix):

@GodGirl360 Look back at the question: At what point, besides the common starting location of the port, do the paths of the two boats cross?

OpenStudy (mathmath333):

yes ^

Directrix (directrix):

>>>They made it clear they didn't want #1. That is why I said it didn't make sense. You are correct.

OpenStudy (mathmath333):

if u dont wanna graph u can solve the two equations http://www.wolframalpha.com/input/?i=y%3D%28-1%2F20%29x%5E2%2B5%2C10y%3D2x%2B20

OpenStudy (anonymous):

Ok, thank you.

Directrix (directrix):

We are happy to help.